OpenStudy (anonymous):

One solution to a quadratic equation is 4 - 5i. Find the equation.

OpenStudy (mertsj):

If 4-5i is a solution then the other solution is 4+5i

OpenStudy (anonymous):

One solution to a quadratic equation is 4 - 5i. Find the equation.

OpenStudy (mertsj):

If you know the two solutions, you can write the factors. Multiply the factors together. Set the product equal to 0.

OpenStudy (anonymous):

(x - 4)2 = 36

OpenStudy (mertsj):

\[[x-(4-5i)][x-(4+5i)]=0\] \[x^2-x(4+5i)-x(4-5i)+(4-5i)(4+5i)=0\] \[x^2-4x-5ix-4x+5ix+16-25i^2=0\] \[x^2-8x+16+25=0\] \[x^2-8x+41=0\]